# Positive Effects of Financing

### v3

@@ Abstract
Today’s decisions on climate change mitigation affect the damage that future generations will bear. Discounting future benefits and costs of climate change mitigation is one of the most critical components of assessing efficient climate mitigation pathways. We extend the DICE model with stochastic discount rates to reflect the uncertain nature of discount rates. Stochastic rates give rise to a stochastic mitigation strategy, resulting in all model quantities becoming stochastic.
We show that the optimization procedure of the DICE model induces intergenerational inequality: lacking a mechanism to regulate burden, future generations have to bear higher costs from abatement and damage relative to GDP.
Further, we show that considering uncertainty of discount rates and their feedback to abatement policies, which can be interpreted as successive re-calculation, increases intergenerational inequality (and adds additional risks). Motivated by this, we consider additional financing risks by investigating two modifications of DICE. We find that allowing financing of abatement costs and considering non-linear financing effects for large damages improves intergenerational effort sharing. To conclude our discussion of options to improve intergenerational equity in an IAM, we propose a modified optimization to keep costs below 3% of GDP, resulting in more equal distribution of efforts between generations.
@@ 1 Introduction
Climate change is one of the most significant risks of the next centuries [1], as demonstrated by its manifold impacts on society [2]. Due to the persistent nature - on human time scales - of the main greenhouse gas carbon dioxide [3], a transition to (net) zero emissions is necessary to limit global warming to any constant level. A diverse set of approaches have been developed to find the optimal pathway towards such an emission-neutral world. One well-established method for the assessment of transition pathways are integrated assessment models (IAMs) of the climate and economic systems; for an overview of the developments on IAMs see [4, 5].
@@
We focus on the DICE model [6, 7] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [8, 9, 10, 11, 12]. While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points [13, 14, 15, 16, 17, 18, 19], natural feedback processes such as permafrost [9] or abstract catastrophic risks [20, 21]. Since most damages of climate change will occur over long time scales, value-based decisions on the discount rate and optimization function have a strong influence on the resulting optimal pathways [22, 23, 24, 25, 26]. Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks [13] or inequalities in the distribution of damages [27]. Most of these contributions aim to provide an estimate of the societal costs of emitting one additional ton of carbon dioxide, the so-called social cost of carbon (SCC), which is an important determinant for benefit-cost analysis of climate mitigation. Model extensions updating IAMs [11, 12, 19, 28] show a wide range of potential SCC, while also econometrics-based SCC estimates for single impacts such as mortality [29] or energy consumption [30], but also general SCC [31] emphasize the need to improve the precision of the estimated SCC, especially by including, so far ignored, additional impact channels. Discount rates, in particular, are an essential component of IAM modeling. Thus, the discussion between proponents of a descriptive approach based on observation of market returns like in DICE [7] versus a normative approach to discounting as in the Stern review [22] continues. The assumptions on discounting have already been discussed in the context of intergenerational [32] and intragenerational inequality [33]. Recent studies [34, 35] show that dynamic modeling of discount rates might allow a more precise assessment of the costs of damage and abatement. While our focus lies on the effect of uncertainty about discount rates and stochastic adaptation of abatement policies on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [36], e.g. carbon taxation, [37].
@@
Parallel to this work, an updated DICE-2023 model was developed and published [38] - one of the most prominent updates includes the consideration of uncertainty in the discount rate through the inclusion of consumption growth uncertainty and riskiness of climate investments. These uncertainties are translated into certainty-equivalent discount rate, such that the overall model remains deterministic. As we show in eq. 11, even a stochastic interest rate without feedback processes within the model is equivalent to a deterministic approach.
Complementary to the potential uncertainties in future consumption growth and riskiness of climate investments, we consider uncertainty about the rate of time preference $\rho$ itself. For this, we use a standard interest-rate model - the Hull-White model [39] to model stochastic fluctuation of the rate of time preference $\rho$, interacting with an abatement policy adjusting to this uncertainty.
1.1 Extensions to the DICE Model
@@
We examine the optimization of the DICE model with respect to uncertainty in the rate of time-preference, and its consequences for intergenerational equity. Measuring the costs per GDP (section 4.4.5) of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations. We use an implementation of the DICE model with a flexible time discretization [40] and couple this model with a comprehensive modeling toolbox for financial markets [41] to enable the integration of valuation methods for complex financial products in IAMs.
@@ 1.2 Interest Rates
We add stochastic interest rates - through modeling the rate of time preference $\rho$ as stochastic - to the model, and introduce a stochastic abatement policy. This policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calculation of the model. Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that influence intergenerational equity: Financing of abatement costs and non-linear discounting of costs, [42]. As abatement is a planned investment process, we assume that these costs may be financed by loans with uncertain interest rates. Furthermore, as large projects and investments often incur additional costs [43, 44], we introduce increasing financing costs for large payments via a non-linear discounting. These higher costs for high damages can be viewed as (additional) financing risks. We provide an overview of the extensions in fig. 1. More background is provided in the methods section.
@@ 2.1 Intergenerational Inequality
To find the optimal mitigation pathway, the standard DICE model equates the increment of the cost of damage $\mathrm{d}C_{\mathrm{D}}(t)$ to the increment of the cost of abatement $\mathrm{d}C_{\mathrm{A}}(t)$, both weighted by the welfare-cost-sensitivity $\mathrm{d}V/\mathrm{d}C$ and the discount factor $\frac{N(0)}{N(t)}$,
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}\mu(s)}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (8)
Considering a simplified 1-parameter abatement model parameterizing the abatement policy by $T^{\mu=1}$, the time by which 100% abatement is achieved, this simplifies to
$\int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ +\ \int_{0}^{T}\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}\ \frac{N(0)}{N(t)}\mathrm{d}t\ \stackrel{{\scriptstyle!}}{{=}}\ 0\text{.}$ (10)
The conditions (8) and (10) are plausible if a single agent optimizes its cost but may be problematic if the costs are distributed between different generations. Note that the two individual contributions in eq. 10 occur at different times. fig. 2 shows the temporal distribution of the weighted functions $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$.
Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$. The optimization is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost, as visible by the shift between the maxima.
Hence, marginal gains of one generation are equated to marginal losses of another generation. The optimization does not equalize the burden but the weighted marginal burden. In this sense, the optimization is indifferent to the temporal distribution of the absolute burden. This may generate intergenerational inequality.
@@ 2.2 Including stochastic interest rates into DICE
We see in Equation 8 that both the welfare-cost-sensitivity $\frac{\mathrm{d}V(t)}{\mathrm{d}C(t)}$ and the discount factor $\frac{N(0)}{N(t)}$ (i.e., the numeraire ratio) are important to determine the inter-temporal optimal pathway.
Including a stochastic interest rate model, the numéraire $N$ becomes stochastic. Allowing for (stochastic) adaptation of the abatement policy $\mu$ to the interest rate level, all model quantities become stochastic.
We use a classical Hull-White model to model uncertainty in the rate of time preference $\rho$ details are provided in the methods (12). The model generates the same expected discount rate as for the deterministic DICE model. This implies that combining the stochastic interest rates with a deterministic abatement model results in the same emission pathways as for the deterministic model. As is well known, from the literature, the level of the discount rate has a large influence on optimal abatement paths - we show in the supplementary information, that an increasing interest rate level defers abatement, while increasing volatility speeds up abatement, as we further discuss in [40].
@@
However, allowing for a stochastic abatement policy that adapts to the interest rate level, e.g., under the model (13) presented in the methods, we see an adverse effect: the expected abatement is smaller, the expected emissions are higher and the expected total cost increase significantly for later years. Note that restricting the stochastic abatement model (13) to the case $a_{1}=0$ agrees with the abatement model (5). The effect is hence solely created by allowing the change of abatement speed by interest rate change $a_{1}$ to be a free parameter.
@@
fig. 3 depicts the results of stochastic interest rates, both in expectation and in a percentile measure (VaR), on emissions, optimal abatement, damage and total cost-per-GDP. These quantities are depicted for the classical deterministic DICE model (grey) and for the DICE model with stochastic interest rates and stochastic abatement policy (green), where we also show the 10th and 90th percentile values.
@@
Figure 3: Classical DICE (grey), Classical DICE with reduce Abatement Model (5) (blue) and Stochastic DICE Model (green). Restricting the abatement policy to the 1-parameter model (5) has a comparably small effect on the emission pathways. Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost). Shaded area shows the 10th to 90th percentiles.
@@
This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles). This is qualitatively independent of the interest rate level.
Such an effect is known under the name of “convexity” in the context of the valuation of financial derivatives: While (11) shows that stochastic interest rates will result in the same cost and emission paths for the deterministic abatement strategy, we do not recover these cost and emissions for the stochastic abatement strategy as cost and emission are convex functions in the interest rates. Note that for a convex function $f$ and a random variable $X$ Jensen’s inequality implies $f(\mathrm{E}(X))\leq\mathrm{E}(f(X))$.
@@
The strength of the effect increases with increasing interest rate volatility. A corresponding analysis for different volatility levels can be found in Variation of the Interest Rate Volatility of the supplementary information. For the limits cast of vanishing interest rate volatility $\sigma=0$, we recover the classical deterministic DICE model.
2.3 Including funding of abatement and funding risks of damage into DICE
As shown in fig. 3, costs are distributed unevenly across generations, with higher total costs relative to GDP for generations in the far future.
We investigate two model extensions that alter the distribution of cost and thus will influence intergenerational equity: First, we modify the abatement costs by allowing these to be funded over time, resulting in interest that has to be paid. Second, we include a non-linear dependency of the discount factor for the damage cost on the magnitude of the cost. This can be viewed as the additional cost associated with financing risk of very large amounts [42]. Details on this extension and parameters are provided in the methods section 4.4.4.
@@ 2.3 Including funding of abatement and funding risks of damage into DICE
We find that both extensions improve intergenerational equity in the classical deterministic DICE model, fig. 4, and in the DICE model with stochastic interest rate, fig. 5. In the latter, intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles.
Figure 4: Deterministic model: Financing extensions lead to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost.
Figure 5: Considering additional costs leads to faster abatement compared to standard model a emissions, b abatement, c damages, and d total GDP relative cost. Shaded area shows the 10th to 90th percentiles (VaR).
@@
Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, present generations still face the lowest costs.
Figures 4 and 5 depict the instantaneous cost-per-GDP ($C(t)/GDP(t)$). For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [45] to calculate burdens experienced during each birth-year cohorts projected lifetime.
@@
We measure the burden in percentage-of-GDP (section 4.4.5), where we choose 3% as the neutral level (green), [46]. While differences are small for the time covered by population projections (fig. 6 and fig. 7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century. Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset. The deterministic model keeps the 3%-limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.

### v4

# Intergenerational Equitable Climate Change Mitigation:

@@ Abstract
Climate mitigation decisions today affect future generations, raising questions of intergenerational equity. Integrated assessment models (IAMs) rely on discounting to evaluate long-term policy costs and benefits. Using the DICE model, we quantify how optimal pathways distribute abatement and damage costs across cohorts. Unconstrained optimization creates intergenerational inequality, with future generations bearing higher costs relative to GDP. Extending the model with stochastic discount rates, we show that discount-rate uncertainty significantly amplifies this inequality. We consider two independent extensions: the financing of abatement costs and the modeling of nonlinear financing costs under large damages. Both extensions can materially improve intergenerational equity by distributing mitigation efforts more evenly. As an illustration, we present a modified DICE model whose optimal pathway limits generational costs to 3 % of GDP, leading to more equitable effort sharing. Our proposed model extensions are model-agnostic, applicable across IAMs, and compatible with alternative intergenerational equity metrics.
[…]
$\printcontents[main]$l1
@@ 1 Introduction
Climate change is one of the largest risks of the next centuries [1], as demonstrated by its manifold impacts on society [2]. Owing to the persistent nature of the main greenhouse gas carbon dioxide [3] on human time scales, a transition to (net) zero emissions is necessary to limit global warming to any constant level. A well-established method for the assessment of transition pathways are integrated assessment models (IAMs) of climate and economic systems [4, 5].
@@
We focus on the DICE model [6, 7, 8] as a simple IAM, which is frequently used to demonstrate the impacts of modifications to the original model due to its simplicity [9, 10, 11, 12, 13]. While IAMs were first set up as deterministic, stochastic shocks have been included to consider the risk of tipping points [14, 15, 16, 17, 18, 19, 20], natural feedback processes such as permafrost [10] or abstract catastrophic risks [21, 22]. Because most of the damages from climate change will occur over long time scales, value-based decisions about the discount rate and the optimization function have a strong influence on the resulting optimal paths [23, 24, 25, 26, 27]. Stochastic approaches also require more advanced risk evaluations than Monte Carlo averaging to capture the full extent of tail risks [14] or inequalities in the distribution of damages [28].
Discount rates, in particular, are an essential component of IAM modeling. Thus, the discussion between proponents of a descriptive approach based on observation of market returns as in DICE [7] versus a normative approach to discounting as in the Stern review [23] continues. The assumptions on discounting have already been discussed in the context of intergenerational [29] and intragenerational inequality [30]. Recent studies [31, 32] show that dynamic modeling of discount rates might allow a more precise assessment of the costs of damage and abatement. While our focus lies on the effect of uncertainty about discount rates and stochastic adaptation of abatement policies on intergenerational equity, we acknowledge that intergenerational equity can be tackled with other policy extensions [33], e.g. carbon taxation, [34].
@@
Parallel to this work, an updated DICE-2023 model was developed and published [8] - one of the most prominent updates includes the consideration of uncertainty in the discount rate through the inclusion of consumption growth uncertainty and riskiness of climate investments. These uncertainties are translated into certainty-equivalent discount rate, such that the overall model remains deterministic. As we show in eq.˜9, even a stochastic interest rate without feedback processes within the model is equivalent to a deterministic approach.
In contrast to this deterministic treatment of interest rate risk, we consider uncertainty about the rate of time preference $\rho$ by modeling it as a stochastic process using a standard interest rate model. Introducing adaptation of the abatement policy, all quantities become stochastic.
We then show that the optimal mitigation pathway derived from an IAM may exhibit substantial intergenerational cost imbalances, which are critically exacerbated by uncertainties in interest rates. We propose two novel model extensions that markedly improve cost-based intergenerational equity, thus enhancing the robustness and fairness of modeling global climate mitigation, providing a crucial political impulse for effective climate action. fig.˜1 provides an overview of the extensions.
Here, we interpret intergenerational equity as ensuring that no cohort bears a disproportionate share of climate-related mitigation and damage costs relative to its economic capacity. To allow meaningful comparisons between generations of differing wealth and productivity, we measure costs as a percentage of GDP, following principles in public finance for sustainable debt and investment burdens [35]. As an illustrative benchmark rather than a normative prescription, we adopt a limit of 3 % of GDP for combined mitigation and damage costs, drawing on similar proposals for a “fair share” of national income devoted to climate action and its consistency with typical upper bounds for long-term public investment commitments. We recognise that alternative justice principles, such as egalitarian, prioritarian, or sufficiency-based approaches [29], would lead to different equity metrics and thresholds. Our results are therefore one contribution to the broader debate on embedding intergenerational fairness in integrated assessment models, but the model extensions we propose can be readily combined with such alternative metrics.
{delete}
2 Method: Extensions to the DICE Model
@@
We examine the optimization of the DICE model with respect to uncertainty in the rate of time-preference, and its consequences for intergenerational equity. Measuring the costs per GDP (section˜5.2.2) of the optimal (calibrated) climate emissions path in the standard model, we find that these costs are unequally distributed across generations. We use an implementation of the DICE model with a flexible time discretization [46] and couple this model with a comprehensive modeling toolbox for financial markets [36] to enable the integration of valuation methods for complex financial products in IAMs.
@@ 2.1 Interest Rates
We add stochastic interest rates - through modeling the rate of time preference $\rho$ as stochastic - to the model, and introduce a stochastic abatement policy. This policy mimics the process of reacting to changes in the interest rate level by repeated (over time) re-calculation of the model. Since intergenerational inequality is related to the sensitivity of the utility and the discount factor, we introduce two model extensions related to these two parts that influence intergenerational equity: Financing of abatement costs and non-linear discounting of costs, [47]. As abatement is a planned investment process, we assume that these costs may be financed by loans with uncertain interest rates. Furthermore, as large projects and investments often incur additional costs [hallegatte_why_2007, shane_construction_2009], we introduce increasing financing costs for large payments via a non-linear discounting. These higher costs for high damages can be viewed as (additional) financing risks.
@@ 3.1 Intergenerational Inequality
The optimality condition of the IAM’s emission path balances the sensitivity of the two cost components, abatement cost and damage cost, to policy changes (see eq.˜5). We individually depict their temporal distribution: the weighted marginal change in abatement and damage cost with respect to a policy change in fig.˜2 and find that an unconstrained IAM determines the optimality of the emission path by balancing the near-time abatement cost changes with the future time damage cost changes.
Figure 2: The sensitivity $\frac{\mathrm{d}C_{\mathrm{A}}(t)}{\mathrm{d}T^{\mu=1}}$ and $\frac{\mathrm{d}C_{\mathrm{D}}(t)}{\mathrm{d}T^{\mu=1}}$ of abatement cost $C_{\mathrm{A}}$ and damage cost $C_{\mathrm{D}}$ on changes of the abatement policy $T^{\mu=1}$. The optimization is balancing the two areas under/above these curves - which is ultimately not resulting in an intergenerational balancing of cost. The two sensitivities come with a significant time shift.
This may already constitute intergenerational inequality, as within that optimization the absolute burden of near-time abatement cost may be much lower than future-time damage cost.
As changes in the abatement policy at time $s$ affect the abatement cost at time $s$ but damage cost at all future times $t\geq s$, we define $\gamma_{\mathrm{D}}(s;t)$ as the damage cost in $t$ per abatement cost in $s$ induced by a change in the abatement policy $\mu(s)$, see (6). We show that at the optimal emission path $\gamma_{\mathrm{D}}(s;t)$ is a probability distribution; a proof can be found in Supplement Section B. This allows us to define the expected time-delay between an abatement and its effect on future damage cost, and we are able to determine it numerically. fig.˜3 depicts the density $\gamma_{\mathrm{D}}(s)$.
Figure 3: The density $t\mapsto\gamma_{\mathrm{D}}(s;t)$ for $s=10,20,50$.
We find that in the classical DICE 2016 model the expected time of damage cost per avoided abatement cost is approximately 70 years after the time of abatement. This observation agrees with fig.˜2, indicating that balancing marginal abatement and damage cost is a matter of intergenerational equity. We depict the temporal distribution of the total cost in fig.˜4, fig.˜5.
3.2 Cost per GDP
Since the optimal emission path balances marginal cost changes only but does not consider cost levels, we like to analyze the temporal distribution of the cost. To measure the burden borne by a generation, we consider the total cost per GDP. Putting nominal values relative to the GDP is a common measure, for example, to assess national debt [35]. Here, it reflects that a wealthier or more productive generation can carry a greater burden, mimicking an effort-sharing scheme, and also considering the greater capacity of future generations under the expected baseline growth in DICE.
@@ 3.3 Including stochastic interest rates into DICE
Discounting plays a fundamental role in the linkage of present and future cost. We assess how uncertainty in discount rates alters the distribution of cost in the optimal emission paths.
We use a classical Hull-White model to model uncertainty in the rate of time preference $\rho$.
If interest rates are stochastic but the utility function remains deterministic, adding stochastic interest rates does not change the interaction with the IAM, and hence the optimal emission path. However, this changes if one assumes that the abatement policy is adapted to the (stochastic) changes in the interest rate level.
We allow for a stochastic abatement policy that adapts to the interest rate level, using the parametric model (10). This reflects the possibility that the abatement policy can be adjusted to the interest rate scenarios.
@@
With the stochastic abatement policy, we see an adverse effect: the expected abatement is smaller, the expected emissions are higher, and the expected total cost increases significantly for later years. Since by restricting a parameter of the stochastic abatement policy ($a_{1}=0$ in (10)) it agrees with the deterministic abatement model (2), it is apparent that the effect is created solely by allowing the abatement speed to depend on the interest rate level (i.e., $a_{1}$ being a free parameter in (10)).
@@
fig.˜4 depicts the results of stochastic interest rates, both in expectation and in a percentile measure (VaR), on emissions, optimal abatement, damage, and total cost-per-GDP.
@@
Figure 4: Classical DICE (grey), Classical DICE with reduced Abatement Model (2) (blue) and Stochastic DICE Model (red). Restricting the abatement policy to the 1-parameter model (2) has a comparably small effect on the emission pathways. Allowing adaptation of the abatement policy to stochastic interest rates increases intergenerational inequality and exhibits a significant risk of even higher intergenerational inequality (a: emissions, b: abatement, c: damages, and d: total GDP relative cost). Shaded area shows the 10th to 90th percentiles. The deterministic model (blue) uses the certainty-equivalent discount rate $\bar{r}$ from (9); therefore, differences in the red line arise solely from interest-rate–driven policy adaptation ($a_{1}\neq 0$) and damage nonlinearity, not from a change in average discounting.
@@
This shows that under the stochastic interest rate and stochastic abatement model, the optimal policy will increase intergenerational inequality and, in addition, exhibit the risk of further increases in intergenerational inequality (as given by the percentiles). In other words, the DICE model exhibits convexity with respect to the interest rates.
@@
This effect is qualitatively independent of the interest rate level. The strength of the effect increases with an increase in interest rate volatility, LABEL:fig:pathways_stochastic_variation. For the limit case of vanishing interest rate volatility $\sigma=0$, we recover the classical deterministic DICE model.
3.4 Improving Intergenerational Equity
We investigate the effect of the two model extensions “funding of abatement cost” (Section 5.3.3) and “non-linear discounting of damage cost” (Section 5.3.4) on the distribution of cost in the deterministic and stochastic model, i.e., how they improve cost-related intergenerational equity.
The parameter choices are exemplary; other experiments with similar results can be reproduced in the published source code.
3.4.1 Funding of Abatement
We allow funding of abatement costs to reduce the time lag between abatement cost and damage cost. As we account for interest for the funding, this does not alter the effect of discounting but brings the two components to similar economical regions.
Shifting cost via a loan to the future may be considered adverse to intergenerational equity; however, our numerical analysis indicates that it is beneficial as it incentivizes higher abatement.
In a numerical experiment, with a funding period of 60 years with equal-value annual repayment (including interest), funding of abatement reduces the cost in the very far future without substantial increase in the near future (fig.˜5, left). An even more notable contribution is that it significantly reduces the cost risk if discount rates are stochastic (fig.˜5, right).
However, with respect to intergenerational equity, improvements are somewhat unbalanced as a funding of abatement still leaves a high burden for near future generations.
3.4.2 Funding Risks of Damage
We add a non-linear discounting of damages. This model extension allows the discount factor to depend on the magnitude of the cash flow, mimicking that funding of larger cash flows requires a larger premium to compensate for default risk or other frictions, a default compensator. As such, larger damages get a larger weight; consequently, the abatement policy optimization will prefer emissions paths that avoid very large damages, which is beneficial with respect to intergenerational equity.
@@ 3.4.2 Funding Risks of Damage
In our numerical experiment, we consider a significant default compensator for damage cost over 3% of the GDP, effectively limiting cost of optimal emission paths to $3\%$. We then find strong improvements to intergenerational equity in the deterministic DICE model and the DICE model with a stochastic interest rate, as shown in fig.˜5. Intergenerational equity is improved with respect to the expectation and the uncertainty in the respective optimal emission pathway, represented by the 10th and 90th percentiles. The associated model quantities (emission, abatement and damage) are depicted in LABEL:fig:pathways_deterministic_with_extension and LABEL:fig:pathways_stochastic_with_extension.
Figure 5: The cost-per-GDP in a the model with deterministic interest rates and b the model with stochastic interest rates, for the otherwise unmodified model (red) and the model with the extension “funding of abatement cost” (blue) and “GDP limited non-linear discounting” (green). Shaded area shows the 10th to 90th percentiles.
@@
Due to the limits of scaling up abatement immediately and the inherent preference for current generations in the DICE model, the present generations still face the lowest costs.
@@
For an assessment of the generational burdens we consider the lifetime-average value using population projections from 2015 to 2100 [37] to calculate burdens experienced during each birth-year cohorts projected lifetime. We measure the burden in percentage-of-GDP (section˜5.2.2), where we choose 3% as the neutral level (green), [38]. While differences are small for the time covered by population projections (fig.˜6 and fig.˜7), extrapolation using the fixed 2100 life expectancy shows a clear divergence at the end of the century. Notably, the stochastic variant does not stay exactly below 3% of GDP on average since high-risk scenarios are not fully offset. The deterministic model keeps the 3%-limit and illustrates the equal burden between generations enabled by imposing GDP relative cost limits.